suppose of production possibilities frontier include the following combinations

cars washing machine
0 1000
100 600
200 0


Q.1 what is the cost of producing nd additional car when 150 cars are being produced?
Q/2 WHAT IS THE COST OF PRODUCING ND ADDITIONAL WASHING MACHINE WHEN 50 CARS ARE BEING PRODUCED? WHEN 150 CARS ARE BEING PRODUCED?
1
Expert's answer
2015-09-09T00:00:43-0400
Q.1 The cost of producing an additional car is (600 - 0)/(200 - 100) = 6 machines when 150 cars are being produced.
Q.2 The cost of producing an additional washing machine is (100 - 0)/(1000 - 600) = 0.25 car, when 50 cars are being produced.
The cost of producing an additional washing machine is (200 - 100)/(600 - 0) = 1/6 car when 150 cars are being produced.